Using the following chemical equation how many grams of hydrochloric acid (HCl) are consumed reacting with 20.43 g of calcium hydroxide?

2HCl + Ca(OH)2 -> 2H2O + CaCl2

1
Expert's answer
2021-02-15T02:47:52-0500

Solution:

The balanced chemical equation:

Ca(OH)2 + 2HCl → CaCl2 + 2H2O

According to the equation above: n(Ca(OH)2) = n(HCl)/2


Moles of Ca(OH)2 = Mass of Ca(OH)2 / Molar mass of Ca(OH)2

The molar mass of Ca(OH)2 is 74.093 g mol-1.

Hence,

n(Ca(OH)2) = 20.43 g / 74.093 g mol-1 = 0.2757 mol


n(HCl) = 2 × n(Ca(OH)2) = 2 × 0.2757 mol = 0.5514 mol


Moles of HCl = Mass of HCl / Molar mass of HCl

Mass of HCl = Moles of HCl × Molar mass of HCl

The molar mass of HCl is 36.458 g mol-1.

Hence,

m(HCl) = 0.5514 mol × 36.458 g mol-1 = 20.1029 g = 20.10 g


Answer: 20.1 grams of hydrochloric acid (HCl) are consumed.